OpenStudy (anonymous):

100^(X+2)=1000^(X-1)

OpenStudy (anonymous):

\[100^{x+2}=1000^{x-1}\rightarrow 10^{2^{x+2}}=10^{3^{x-1}}\] \[2x+4=3x-3\rightarrow x=7\]